4. Here are some facts about Washington State. Source: State Health Facts 6 pts) e Washington total population: 7,297,300 Number of children (0-18) in Washington State: 1,720,900 Percentage of Washington population that is 65 years or older: 15% What percentage of Washington States population is children (0-18)? a. b. How many people in Washington state are 65 years or older? 7297 3σο χ .IS = lo q4S1SPeople or as-6s landrix College gave the following t arademic
Answer #1

Here we are given in Washington total population =7297300

No. Of children in Washington state=1720900

Percentage of people that is 65 or more year old=15%

Then we are to find

1.percentage of people between 0-18

Ans=(1720900/7297300)*100=23.58%

2.how many people are old.

Ans=(7297300*15)/100

1094595